# Dual-View Visual Contextualization for Web Navigation

### v1

@@ 3.2 Context enhancement
Figure 4: Dual-VCR-enhanced element ranker.. We contextualize the candidate element (denoted by normal-$\star$) with its neighbors in the screenshot, using both the visual features (by [20]) and textual features (extracted from the HTML document). Positional embeddings are added to specify neighbor elements, learning their spatial relationships and pairing the textual features with visual features. This dual-view contextualized representation is used to rank the candidate element, measuring its relevance to the current task.
@@ 3.3 Dual-VCR-enhanced element ranker
Figure 5: Dual-VCR-enhanced action predictor. Given the top-$K$ candidate elements (three in the figure, marked with normal-$\star$), Dual-VCR appends each with its neighbor elements. The resulting HTML snippet, together with the task description and previous actions, is then fed into an LLM for predicting the next action.
@@ 4 Experimental Results
Dataset. We validate Dual-VCR on Mind2Web [8], a comprehensive benchmark for real-world web navigation. Unlike other benchmarks based on simulated websites with only a few HTML elements, Mind2Web uses over 100 real-world websites with thousands of HTML elements. Concretely, they provide over 2K open-ended tasks collected from 137 real-world websites across 31 different domains, including travel, shopping, public service, etc (Table 1). A task consists of a sequence of actions, each comprising a pair of an actionable HTML element (e.g., “[textbox] To”) and an operation (e.g., “Type Toronto”). Mind2Web provides three common operations: Click, Type, and Select. For Type and Select, an additional argument is required (e.g., Select “Queen” for choosing the bed size). In total, Mind2Web provides 1,009 tasks for training split.
@@
Other Methods. A number of studies [16, 36, 23, 32, 2, 26, 18, 31, 30] have explored web navigation but mainly worked on simplified websites [16, 36], which deviate from the focus of our study. Our attention is instead directed towards real-world scenarios involving various real-world websites with extensive raw HTML documents (e.g., Mind2Web). We have identified a few concurrent works [9, 10, 37, 11, 14, 3] exploring Mind2Web, but they mostly focus on (i) large-scale pre-training, requiring substantial amounts of pre-training HTML data, or (ii) evaluating the potential of recent vision-and-language models (e.g., GPT4-V [29]) as a web agent. As their codes or pre-training datasets have not been released yet, replicating their work would be prohibitively costly. We thus do not consider them in our studies.
Our Models. Aligned with MindAct, we use the same DeBERTabase [13] / Flan-T5base [6] for our ranker / action predictor, repsectively. For visual features extraction, we utilize Pix2Struct [20]’s ViT (pre-trained on screenshots) as the visual backbone and apply ROI Align [12] on the element’s region. We use two linear layers to project visual features into textual embedding space.
@@
Training Details. For the Dual-VCR-enhanced element ranker, we learn its projection layer, positional embeddings, and ranking LM while keeping its ViT frozen. The LM outputs each element’s importance score for action prediction through a sigmoid activation function. This score is optimized using a binary cross-entropy loss, where the ground-truth element serves as a positive example, and elements randomly sampled from the webpage are considered negative examples. For the Dual-VCR-enhanced action predictor, we adopt the configuration from MindAct, optimizing its parameters with the language modeling loss. Both our ranker and predictor are trained with a batch size of 32 and a learning rate of 3e-5 for 5 epochs. We utilize a single Nvidia A6000 48GB GPU during training. Please refer to the supplementary materials for more details.
@@ 4.1 Effectinvess of Dual-VCR
Dataset Website # Tasks Avg # HTML
@@
MiniWoB++ [16] Simplified 100 28 500
Mind2Web [8] Real-world 2,350 1,135 44,402
@@
Ranking performance. Table 2 summarizes the ranking results across different top-$K$ candidate elements. First, we see that incorporating the visual neighbor elements’ HTML text (Dual-VCRvnei-txt) consistently and significantly outperforms MindActRank on all Recall@$K$s (e.g., 37.3% vs. 25.4% on Recall@1, 79.3% vs. 73.5% on Recall@10), suggesting that contextualizing the element with its neighbors indeed helps find the target element. Second, the candidate element’s visual features (Dual-VCRvis) leads to the notable improvements over MindActRank (e.g., 70.2% vs. 61.0% on Recall@5). This implies that the visual features offer additional context in differentiating HTML elements, compared to using only its HTML text. Lastly, Dual-VCRvnei-txt+vis achieves a further boost by leveraging both visual neighbors’ HTML text and visual features (e.g., 38.4%/90.1% on Recall@1/@50).
@@ 4.2 Analysis
Effects of the number of visual neighbors. We ablate the impact of varying sizes of visual neighbors, starting with  Table 8, which shows its effect on the ranker while maintaining the same action predictor (MindActPred). We observe a linear correlation between the size of visual neighbors and their ranking/action prediction performance. For instance, increasing the size of neighbors up to five shows consistent improvements (e.g., 89.1$\rightarrow$90.1on Recall@50 and 75.1$\rightarrow$78.6on Op. F1). However, considering too many neighbors (e.g., the size of ten) hurts the performance. For example, increasing the size from five to ten decreases the element accuracy from 46.0% to 45.2%. We also see a similar pattern when ablating the effect of the visual neighbor size on the action predictor (Table 9). Concretely, while keeping the same ranker (Dual-VCRvnei-txt+vis), the action performance increases up to the size of five (e.g., 46.0$\rightarrow$47.0on Ele. Acc) but decreases when the size becomes ten (e.g., 46.2% on Ele. Acc). Overall, this suggests that choosing an appropriate number of neighbors is necessary for both element ranking and action prediction, consistent with Table 6, where using the entire HTML documents hurts.
@@ Acknowledgments
  • [itemsep=5pt,topsep=1.5pt]
  • Appendix A: Model implementation details (cf. §3.3,  §3.4, and §4 of the main text).
@@
Appendix A Model implementation details
@@ Appendix C Additional Experiments
Table 11: Dual-VCR with a larger predictor. We increase the size of predictor from Flan-T5base to Flan-T5large. Even with the larger predictor, Dual-VCR notably outperforms the baseline, showing the complementarity of Dual-VCR and LLMs.

### v2

@@ 3.2 Context enhancement
Figure 4: Dual-VCR-enhanced element ranker.. We contextualize the candidate element (denoted by $\star$) with its neighbors in the screenshot, using both the visual features (by [20]) and textual features (extracted from the HTML document). Positional embeddings are added to specify neighbor elements, learning their spatial relationships and pairing the textual features with visual features. This dual-view contextualized representation is used to rank the candidate element, measuring its relevance to the current task.
@@ 3.3 Dual-VCR-enhanced element ranker
Figure 5: Dual-VCR-enhanced action predictor. Given the top-$K$ candidate elements (three in the figure, marked with $\star$), Dual-VCR appends each with its neighbor elements. The resulting HTML snippet, together with the task description and previous actions, is then fed into an LLM for predicting the next action.
@@ 4 Experimental Results
Dataset. We validate Dual-VCR on Mind2Web [8], a comprehensive benchmark for real-world web navigation. Unlike other benchmarks based on simulated websites with only a few HTML elements, Mind2Web uses over 100 real-world websites with thousands of HTML elements. Concretely, they provide over 2K open-ended tasks collected from 137 real-world websites across 31 different domains, including travel, shopping, public service, etc (Table 1). Please see more details in the supplementary material.
@@
Our Models. Aligned with MindAct, we use the same DeBERTabase [13] / Flan-T5base [6] for our ranker / action predictor, repsectively. For visual features extraction, we utilize Pix2Struct [20]’s ViT (pre-trained on screenshots) as the visual backbone and apply ROI Align [12] on the element’s region. We use two linear layers to project visual features into textual embedding space. Please see the supplementary materials for details on the model training.
@@ 4.1 Effectinvess of Dual-VCR
Dataset # Websites Website # Tasks Avg # HTML
@@
MiniWoB++ [16] 100 Simplified 100 28 500
Mind2Web [8] 137 Real-world 2,350 1,135 44,402
@@
Ranking performance. Table 2 summarizes the ranking results across different top-$K$ candidate elements. First, we see that incorporating the visual neighbor elements’ HTML text (Dual-VCRvnei-txt) consistently and significantly outperforms MindActRank on all Recall@$K$s (e.g., 37.3% vs. 25.4% on Recall@1, 79.3% vs. 73.5% on Recall@10), suggesting that contextualizing the element with its neighbors indeed helps find the target element. Second, the candidate element’s visual features (Dual-VCRvis) lead to notable improvements over MindActRank (e.g., 70.2% vs. 61.0% on Recall@5). This implies that the visual features offer additional context in differentiating HTML elements, compared to using only its HTML text. Lastly, Dual-VCRvnei-txt+vis achieves a further boost by leveraging both visual neighbors’ HTML text and visual features (e.g., 38.4%/90.1% on Recall@1/@50).
@@ 4.2 Analysis
Effects of the number of visual neighbors. We ablate the impact of varying sizes of visual neighbors, starting with  Table 8, which shows its effect on the ranker while maintaining the same action predictor (MindActPred). We observe a linear correlation between the size of visual neighbors and their ranking/action prediction performance. For instance, increasing the size of neighbors up to five shows consistent improvements (e.g., 89.1$\rightarrow$90.1on Recall@50 and 75.1$\rightarrow$78.6on Op. F1). However, considering too many neighbors (e.g., the size of ten) hurts the performance. For example, increasing the size from five to ten decreases the element accuracy from 46.0% to 45.2%. We also see a similar pattern when ablating the effect of the visual neighbor size on the action predictor (Table 9). Concretely, while keeping the same ranker (Dual-VCRvnei-txt+vis), the action performance increases up to the size of five (e.g., 46.0$\rightarrow$47.0on Ele. Acc) but decreases when the size becomes ten (e.g., 46.2% on Ele. Acc). Overall, this suggests that choosing an appropriate number of neighbors is necessary for both element ranking and action prediction.
@@ Acknowledgments
  • Appendix A: Model implementation & training details (cf. §3.3,  §3.4, and §4 of the main text).
@@
Appendix A Model implementation & training details
@@ Appendix B Dataset Details
Task Details. The Mind2Web task consists of a sequence of actions, each comprising a pair of an actionable HTML element (e.g., “[textbox] To”) and an operation (e.g., “Type Toronto”). Mind2Web provides three common operations: Click, Type, and Select. For Type and Select operations, an additional argument (e.g., “Toronto”) is required.
@@ Appendix C Additional Experiments
Table 11: Dual-VCR with a larger predictor. We increase the size of the predictor from Flan-T5base to Flan-T5large. Even with the larger predictor, Dual-VCR notably outperforms the baseline, showing the complementarity of Dual-VCR and LLMs.
@@
Existing/Concurrent Works. A number of previous studies [16, 36, 23, 32, 2, 26, 18, 31, 30] have explored web navigation but mainly worked on simplified websites [16, 36], which deviate from the focus of our study. Our attention is instead directed towards real-world scenarios involving various real-world websites with extensive raw HTML documents (e.g., Mind2Web). We have identified a few concurrent works [9, 10, 37, 11, 14, 3] exploring Mind2Web, but they mostly focus on (i) large-scale pre-training, requiring substantial amounts of pre-training HTML data, or (ii) evaluating the potential of recent vision-and-language models (e.g., GPT4-V [29]) as a web agent. As their codes or pre-training datasets have not been released yet, replicating their work would be prohibitively costly. We thus do not consider them in our studies.